What is the y-intercept of 9x-6y=12
Answer by nerdybill(7384)

What is the y-intercept of 9x-6y=12
.
Manipulate the equation into the "slop intercept" form:
9x-6y=12
-6y= -9x+12
y = (9/6)x-(12/6)
y = (3/2)x - 2
.
y-intercept:
(0, -2)